OpenStudy (anonymous):

Which o f the following points lies on the graph of the function y= 4^x (1,0) (4,1) (2,16) (-1,.05)

OpenStudy (anonymous):

(2,16) , plug x=2 and u get correspondingly y=16.

OpenStudy (anonymous):

\[4^1 = 4\]\[4^4=256\]\[4^2 = 16\]\[4^{-1} =\frac{1}{4}\] so it has to be (2,16)

OpenStudy (anonymous):

since comupters were invented, computational speed has multipilied by a factor of 4 about every 3 years. if a typical computer operates with a computational speed s today, write an expression for the speed at which you can expect an equivalent computer to operate after x three year periods